### v2

@@ Appendix A Extended Data
| ResNet50 | 0.869 (0.805-0.926) | 0.843 (0.778-0.905) | 0.945 (0.905-0.977) |
| Phikon | 0.968 (0.928-0.996) | 0.967 (0.932-0.993) | 0.990 (0.974-1.000) |
| Ctranspath | 0.922 (0.862-0.973) | 0.924 (0.874-0.970) | 0.974 (0.949-0.994) |
| UNI | 0.931 (0.872-0.978) | 0.941 (0.890-0.976) | 0.996 (0.989-1.000) |
| CONCH | 0.944 (0.891-0.984) | 0.949 (0.909-0.984) | 0.995 (0.987-0.999) |
| PLIP | 0.955 (0.911-0.988) | 0.943 (0.900-0.982) | 0.974 (0.931-1.000) |
| GPFM | 0.960 (0.915-996) | 0.967 (0.933-0.993) | 0.998 (0.995-1.000) |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.685 (0.577-0.783) | 0.675 (0.577-0.771) | 0.919 (0.875-0.954) |
| Phikon | 0.764 (0.681-0.852) | 0.756 (0.669-0.834) | 0.966 (0.944-0.985) |
| Ctranspath | 0.763 (0.670-0.848) | 0.760 (0.663-0.854) | 0.960 (0.936-0.979) |
| UNI | 0.815 (0.731-0.900) | 0.823 (0.713-0.910) | 0.969 (0.943-0.990) |
| CONCH | 0.774 (0.694-0.862) | 0.760 (0.666-0.864) | 0.957 (0.930-0.979) |
| PLIP | 0.775 (0.690-0.861) | 0.755 (0.646-0.849) | 0.946 (0.915-0.973) |
| GPFM | 0.796 (0.712-0.876) | 0.799 (0.708-0.881) | 0.972 (0.952-0.988) |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.693 (0.618-0.772) | 0.675 (0.602-0.748) | 0.959 (0.944-0.971) |
| Phikon | 0.834 (0.772-0.892) | 0.842 (0.782-0.897) | 0.990 (0.986-995) |
| Ctranspath | 0.818 (0.749-0.884) | 0.830 (0.763-0.882) | 0.985 (0.976-0.993) |
| UNI | 0.897 (0.842-0.946) | 0.898 (0.843-0.937) | 0.996 (0.992-0.999) |
| CONCH | 0.909 (0.854-0.951) | 0.910 (0.860-0.947) | 0.992 (0.988-0.997) |
| PLIP | 0.758 (0.680-0.828) | 0.757 (0.686-0.823) | 0.977 (0.965-0.986) |
| GPFM | 0.899 (0.843-0.945) | 0.890 (0.829-0.936) | 0.996 (0.994-0.999) |
| Models | C-Index |
| ResNet50 | 0.648 (0.543-0.752) |
| Phikon | 0.720 (0.609-0.830) |
| Ctranspath | 0.683 (0.607-0.758) |
| UNI | 0.698 (0.594-0.802) |
| CONCH | 0.667 (0.601-0.733) |
| PLIP | 0.671 (0.602-0.741) |
| GPFM | 0.726 (0.634-0.818) |
| BRCA | BLCA | KIRC | |
| ResNet50 | 0.678 (0.660-0.696) | 0.696 (0.653-0.739) | 0.634 (0.565-0.703) |
| Phikon | 0.770 (0.750-0.790) | 0.692 (0.657-0.727) | 0.708 (0.673-0.743) |
| Ctranspath | 0.704 (0.682-0.726) | 0.670 (0.641-0.699) | 0.671 (0.608-0.734) |
| UNI | 0.722 (0.710-0.734) | 0.690 (0.645-0.735) | 0.674 (0.627-0.721) |
| CONCH | 0.677 (0.636-0.718) | 0.644 (0.544-0.744) | 0.659 (0.583-0.735) |
| PLIP | 0.691 (0.666-0.716) | 0.665 (0.594-0.736) | 0.652 (0.572-0.732) |
| GPFM | 0.757 (0.737-0.777) | 0.711 (0.682-0.740) | 0.697 (0.658-0.736) |
| KIRP | STAD | CESC | |
| ResNet50 | 0.630 (0.562-0.697) | 0.636 (0.572-0.701) | 0.611 (0.517-0.705) |
| Phikon | 0.680 (0.615-0.745) | 0.705 (0.660-0.750) | 0.747 (0.702-0.792) |
| Ctranspath | 0.649 (0.567-0.731) | 0.669 (0.610-0.728) | 0.702 (0.569-0.835) |
| UNI | 0.659 (0.563-0.755) | 0.688 (0.641-0.735) | 0.735 (0.649-0.821) |
| CONCH | 0.623 (0.527-0.719) | 0.636 (0.554-0.718) | 0.677 (0.610-0.744) |
| PLIP | 0.674 (0.631-0.717) | 0.657 (0.584-0.730) | 0.694 (0.588-0.800) |
| GPFM | 0.688 (0.639-0.737) | 0.694 (0.645-0.743) | 0.754 (0.646-0.862) |
| LUAD | CPTAC-LUAD | LUSC | |
| ResNet50 | 0.646 (0.601-0.691) | 0.517 (0.431-0.603) | 0.667 (0.623-0.711) |
| Phikon | 0.738 (0.644-0.832) | 0.577 (0.424-0.730) | 0.779 (0.729-0.829) |
| Ctranspath | 0.633 (0.584-0.682) | 0.624 (0.565-0.683) | 0.703 (0.590-0.817) |
| UNI | 0.673 (0.614-0.732) | 0.575 (0.495-0.655) | 0.746 (0.665-0.828) |
| CONCH | 0.641 (0.568-0.714) | 0.649 (0.608-0.690) | 0.661 (0.586-0.737) |
| PLIP | 0.628 (0.536-0.720) | 0.608 (0.508-0.708) | 0.686 (0.647-0.725) |
| GPFM | 0.739 (0.690-0.788) | 0.626 (0.567-0.685) | 0.777 (0.731-0.822) |
| COADREAD | GBMLGG | SKCM | |
| ResNet50 | 0.632 (0.579-0.685) | 0.724 (0.687-0.761) | 0.699 (0.621-0.777) |
| Phikon | 0.707 (0.646-0.768) | 0.760 (0.740-0.780) | 0.773 (0.732-0.814) |
| Ctranspath | 0.680 (0.633-0.727) | 0.745 (0.716-0.774) | 0.743 (0.710-0.776) |
| UNI | 0.691 (0.644-0.738) | 0.778 (0.754-0.802) | 0.747 (0.714-0.780) |
| CONCH | 0.679 (0.616-0.742) | 0.735 (0.708-0.762) | 0.721 (0.627-0.815) |
| PLIP | 0.653 (0.573-0.733) | 0.738 (0.691-0.785) | 0.709 (0.652-0.766) |
| GPFM | 0.708 (0.663-0.753) | 0.785 (0.756-0.814) | 0.777 (0.716-0.838) |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.757 (0.451-1.000) | 0.749 (0.422-1.000) | 0.927 (0.766-1.000) |
| Phikon | 0.824 (0.501-1.000) | 0.816 (0.477-1.000) | 0.948 (0.809-1.000) |
| Ctranspath | 0.799 (0.458-1.000) | 0.794 (0.443-1.000) | 0.947 (0.822-1.000) |
| UNI | 0.845 (0.569-1.000) | 0.840 (0.552-1.000) | 0.953 (0.824-1.000) |
| CONCH | 0.823 (0.578-1.000) | 0.819 (0.566-1.000) | 0.951 (0.837-1.000) |
| PLIP | 0.762 (0.476-1.000) | 0.753 (0.441-1.000) | 0.928 (0.759-1.000) |
| GPFM | 0.853 (0.559-1.000) | 0.848 (0.540-1.000) | 0.955 (0.832-1.000) |
@@
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.865 (0.788-0.932) | 0.856 (0.776-0.928) | 0.977 (0.958-0.992) |
| Phikon | 0.918 (0.845-0.971) | 0.915 (0.842-0.965) | 0.988 (0.975-0.998) |
| Ctranspath | 0.927 (0.865-0.975) | 0.919 (0.861-0.965) | 0.998 (0.993-1.000) |
| UNI | 0.960 (0.915-1.000) | 0.966 (0.911-1.000) | 1.000 (0.999-1.000) |
| CONCH | 0.934 (0.879-0.981) | 0.933 (0.885-0.986) | 0.996 (0.988-1.000) |
| PLIP | 0.799 (0.714-0.871) | 0.791 (0.698-0.880) | 0.959 (0.926-0.981) |
| GPFM | 0.963 (0.919-1.000) | 0.965 (0.915-1.000) | 0.998 (0.994-1.000) |
@@
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.813 (0.809-0.818) | 0.843 (0.839-0.847) | 0.946 (0.944-0.948) |
| Phikon | 0.893 (0.889-0.896) | 0.901 (0.897-0.904) | 0.975 (0.974-0.977) |
| Ctranspath | 0.857 (0.852-0.860) | 0.880 (0.876-0.883) | 0.965 (0.963-0.967) |
| UNI | 0.897 (0.893-0.900) | 0.905 (0.902-0.908) | 0.977 (0.976-0.979) |
| CONCH | 0.866 (0.862-0.870) | 0.889 (0.885-0.892) | 0.971 (0.969-0.973) |
| PLIP | 0.810 (0.805-0.815) | 0.843 (0.838-0.847) | 0.949 (0.947-0.951) |
| GPFM | 0.894 (0.890-0.897) | 0.908 (0.904-0.911) | 0.978 (0.977-0.979) |
@@
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.937 (0.923-0.950) | 0.938 (0.925-0.951) | 0.986 (0.981-0.990) |
| Phikon | 0.973 (0.964-0.981) | 0.973 (0.965-0.982) | 0.997 (0.996-0.998) |
| Ctranspath | 0.962 (0.952-0.972) | 0.961 (0.951-0.971) | 0.995 (0.992-0.997) |
| UNI | 0.977 (0.967-0.984) | 0.976 (0.968-0.984) | 0.998 (0.997-0.999) |
| CONCH | 0.950 (0.935-0.961) | 0.952 (0.941-0.963) | 0.991 (0.986-0.994) |
| PLIP | 0.943 (0.929-0.954) | 0.940 (0.927-0.951) | 0.989 (0.986-0.993) |
| GPFM | 0.974 (0.965-0.984) | 0.976 (0.968-0.984) | 0.998 (0.997-0.999) |
| Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | 0.725 (0.704-0.746) | 0.735 (0.715-0.757) | 0.930 (0.921-0.938) |
| Phikon | 0.782 (0.762-0.804) | 0.784 (0.763-0.803) | 0.952 (0.945-0.958) |
| Ctranspath | 0.772 (0.752-0.793) | 0.779 (0.757-0.798) | 0.950 (0.943-0.957) |
| UNI | 0.790 (0.770-0.809) | 0.789 (0.770-0.809) | 0.952 (0.945-0.958) |
| CONCH | 0.759 (0.738-0.778) | 0.762 (0.743-0.783) | 0.942 (0.934-0.948) |
| PLIP | 0.747 (0.724-0.768) | 0.755 (0.735-0.775) | 0.941 (0.935-0.949) |
| GPFM | 0.797 (0.776-0.817) | 0.803 (0.784-0.821) | 0.956 (0.950-0.963) |
@@
| ResNet50 | 28.17%(26.63%-29.70%) | 86.52%(85.43%-87.61%) | 57.35% |
| Phikon | 30.78%(29.28%-32.29%) | 87.20%(86.13%-88.27%) | 58.99% |
| Ctranspath | 31.11%(29.58%-32.65%) | 87.51%(86.44%-88.58%) | 60.32% |
| UNI | 33.85%(32.28%-35.42%) | 88.69%(87.64%-89.74%) | 61.27% |
| CONCH | 37.08%(35.40%-38.77%) | 88.51%(87.49%-89.53%) | 62.80% |
| PLIP | 30.83%(29.29%-32.37%) | 88.02%(86.94%-89.09%) | 59.43% |
| GPFM | 34.26%(32.67%-35.84%) | 88.41%(87.32%-89.49%) | 61.33% |
@@
| GPFM (our) | 33 Pubic datasets | 72K | 190M | ViT-L | 307M | UDK |

### v3

@@ Appendix A Extended Data
| ResNet50 | 0.745±0.160 | 0.732±0.174 | 0.906±0.096 |
| Phikon | 0.820±0.155 | 0.810±0.162 | 0.934±0.075 |
| Ctranspath | 0.797±0.161 | 0.792±0.165 | 0.929±0.076 |
| UNI | 0.851±0.133 | 0.848±0.137 | 0.939±0.072 |
| CONCH | 0.820±0.130 | 0.818±0.131 | 0.932±0.077 |
| PLIP | 0.756±0.150 | 0.743±0.164 | 0.899±0.116 |
| CHIEF | 0.790±0.150 | 0.785±0.154 | 0.926±0.082 |
| Prov-Gigapath | 0.856±0.132 | 0.856±0.135 | 0.944±0.065 |
| GPFM | 0.866±0.136 | 0.865±0.142 | 0.946±0.066 |
@@
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | BACH | 0.865 (0.788-0.932) | 0.856 (0.776-0.928) | 0.977 (0.958-0.992) |
| Phikon | BACH | 0.918 (0.845-0.971) | 0.915 (0.842-0.965) | 0.988 (0.975-0.998) |
| Ctranspath | BACH | 0.927 (0.865-0.975) | 0.919 (0.861-0.965) | 0.998 (0.993-1.000) |
| UNI | BACH | 0.960 (0.915-1.000) | 0.966 (0.911-1.000) | 1.000 (0.999-1.000) |
| CONCH | BACH | 0.934 (0.879-0.981) | 0.933 (0.885-0.986) | 0.996 (0.988-1.000) |
| PLIP | BACH | 0.799 (0.714-0.871) | 0.791 (0.698-0.880) | 0.959 (0.926-0.981) |
| CHIEF | BACH | 0.925 (0.865-0.975) | 0.924 (0.859-0.975) | 0.998 (0.993-1.000) |
| Prov-Gigapath | BACH | 0.925 (0.866-0.975) | 0.924 (0.859-0.974) | 0.995 (0.986-1.000) |
| GPFM | BACH | 0.963 (0.919-1.000) | 0.965 (0.915-1.000) | 0.998 (0.994-1.000) |
| ResNet50 | BreakHis | 0.937 (0.923-0.950) | 0.938 (0.925-0.951) | 0.986 (0.981-0.990) |
| Phikon | BreakHis | 0.973 (0.964-0.981) | 0.973 (0.965-0.982) | 0.997 (0.996-0.998) |
| Ctranspath | BreakHis | 0.962 (0.952-0.972) | 0.961 (0.951-0.971) | 0.995 (0.992-0.997) |
| UNI | BreakHis | 0.977 (0.967-0.984) | 0.976 (0.968-0.984) | 0.998 (0.997-0.999) |
| CONCH | BreakHis | 0.950 (0.935-0.961) | 0.952 (0.941-0.963) | 0.991 (0.986-0.994) |
| PLIP | BreakHis | 0.943 (0.929-0.954) | 0.940 (0.927-0.951) | 0.989 (0.986-0.993) |
| CHIEF | BreakHis | 0.961 (0.950-0.972) | 0.961 (0.950-0.971) | 0.995 (0.993-0.997) |
| Prov-Gigapath | BreakHis | 0.974 (0.966-0.983) | 0.974 (0.965-0.982) | 0.998 (0.997-0.999) |
| GPFM | BreakHis | 0.974 (0.965-0.984) | 0.976 (0.968-0.984) | 0.998 (0.997-0.999) |
@@
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | PanCancer-TIL | 0.813 (0.809-0.818) | 0.843 (0.839-0.847) | 0.946 (0.944-0.948) |
| Phikon | PanCancer-TIL | 0.893 (0.889-0.896) | 0.901 (0.897-0.904) | 0.975 (0.974-0.977) |
| Ctranspath | PanCancer-TIL | 0.857 (0.852-0.860) | 0.880 (0.876-0.883) | 0.965 (0.963-0.967) |
| UNI | PanCancer-TIL | 0.897 (0.893-0.900) | 0.905 (0.902-0.908) | 0.977 (0.976-0.979) |
| CONCH | PanCancer-TIL | 0.866 (0.862-0.870) | 0.889 (0.885-0.892) | 0.971 (0.969-0.973) |
| PLIP | PanCancer-TIL | 0.810 (0.805-0.815) | 0.843 (0.838-0.847) | 0.949 (0.947-0.951) |
| CHIEF | PanCancer-TIL | 0.845 (0.841-0.849) | 0.873 (0.869-0.876) | 0.961 (0.960-0.963) |
| Prov-Gigapath | PanCancer-TIL | 0.886 (0.883-0.890) | 0.894 (0.891-0.898) | 0.968 (0.967-0.970) |
| GPFM | PanCancer-TIL | 0.894 (0.890-0.897) | 0.908 (0.904-0.911) | 0.978 (0.977-0.979) |
| ResNet50 | Center-3-TIL* | 0.768 (0.763-0.773) | 0.749 (0.743-0.755) | 0.886 (0.881-0.891) |
| Phikon | Center-3-TIL* | 0.781 (0.776-0.786) | 0.762 (0.756-0.768) | 0.917 (0.913-0.921) |
| Ctranspath | Center-3-TIL* | 0.771 (0.766-0.776) | 0.750 (0.744-0.757) | 0.905 (0.900-0.909) |
| UNI | Center-3-TIL* | 0.917 (0.913-0.921) | 0.914 (0.910-0.918) | 0.929 (0.924-0.932) |
| CONCH | Center-3-TIL* | 0.915 (0.911-0.919) | 0.912 (0.908-0.916) | 0.934 (0.930-0.938) |
| PLIP | Center-3-TIL* | 0.807 (0.802-0.812) | 0.793 (0.788-0.799) | 0.888 (0.884-0.893) |
| CHIEF | Center-3-TIL* | 0.758 (0.753-0.764) | 0.733 (0.727-0.740) | 0.915 (0.911-0.919) |
| Prov-Gigapath | Center-3-TIL* | 0.837 (0.832-0.842) | 0.826 (0.820-0.831) | 0.926 (0.922-0.930) |
| GPFM | Center-3-TIL* | 0.942 (0.939-0.946) | 0.940 (0.937-0.944) | 0.951 (0.948-0.955) |
@@
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | Chaoyang | 0.725 (0.704-0.746) | 0.735 (0.715-0.757) | 0.930 (0.921-0.938) |
| Phikon | Chaoyang | 0.782 (0.762-0.804) | 0.784 (0.763-0.803) | 0.952 (0.945-0.958) |
| Ctranspath | Chaoyang | 0.772 (0.752-0.793) | 0.779 (0.757-0.798) | 0.950 (0.943-0.957) |
| UNI | Chaoyang | 0.790 (0.770-0.809) | 0.789 (0.770-0.809) | 0.952 (0.945-0.958) |
| CONCH | Chaoyang | 0.759 (0.738-0.778) | 0.762 (0.743-0.783) | 0.942 (0.934-0.948) |
| PLIP | Chaoyang | 0.747 (0.724-0.768) | 0.755 (0.735-0.775) | 0.941 (0.935-0.949) |
| CHIEF | Chaoyang | 0.765 (0.743-0.785) | 0.772 (0.749-0.792) | 0.948 (0.942-0.956) |
| Prov-Gigapath | Chaoyang | 0.797 (0.776-0.816) | 0.799 (0.779-0.818) | 0.957 (0.952-0.963) |
| GPFM | Chaoyang | 0.797 (0.776-0.817) | 0.803 (0.784-0.821) | 0.956 (0.950-0.963) |
| ResNet50 | Center-3-Colon* | 0.560 (0.556-0.564) | 0.495 (0.488-0.502) | 0.768 (0.762-0.774) |
| Phikon | Center-3-Colon* | 0.684 (0.679-0.689) | 0.678 (0.671-0.684) | 0.841 (0.835-0.846) |
| Ctranspath | Center-3-Colon* | 0.700 (0.695-0.706) | 0.701 (0.694-0.707) | 0.826 (0.821-0.832) |
| UNI | Center-3-Colon* | 0.724 (0.718-0.729) | 0.724 (0.717-0.730) | 0.868 (0.863-0.873) |
| CONCH | Center-3-Colon* | 0.731 (0.725-0.737) | 0.730 (0.724-0.736) | 0.803 (0.796-0.809) |
| PLIP | Center-3-Colon* | 0.626 (0.621-0.632) | 0.603 (0.596-0.610) | 0.770 (0.763-0.776) |
| CHIEF | Center-3-Colon* | 0.690 (0.684-0.695) | 0.688 (0.682-0.695) | 0.820 (0.814-0.825) |
| Prov-Gigapath | Center-3-Colon* | 0.885 (0.881-0.890) | 0.893 (0.889-0.898) | 0.913 (0.909-0.917) |
| GPFM | Center-3-Colon* | 0.828 (0.823-0.833) | 0.836 (0.831-0.842) | 0.891 (0.886-0.896) |
| Cohort | Balanced ACC | Weighted F1 | AUC | |
| ResNet50 | GasHisDB | 0.953 (0.947-0.958) | 0.954 (0.949-0.959) | 0.992 (0.990-0.993) |
| Phikon | GasHisDB | 0.995 (0.993-0.997) | 0.995 (0.994-0.997) | 1.000 (1.000-1.000) |
| Ctranspath | GasHisDB | 0.980 (0.976-0.983) | 0.980 (0.976-0.983) | 0.998 (0.998-0.999) |
| UNI | GasHisDB | 0.996 (0.994-0.997) | 0.996 (0.994-0.997) | 1.000 (1.000-1.000) |
| CONCH | GasHisDB | 0.981 (0.978-0.985) | 0.981 (0.978-0.985) | 0.998 (0.998-0.999) |
| PLIP | GasHisDB | 0.958 (0.954-0.963) | 0.958 (0.953-0.963) | 0.993 (0.991-0.994) |
| CHIEF | GasHisDB | 0.979 (0.976-0.983) | 0.980 (0.976-0.983) | 0.998 (0.997-0.999) |
| Prov-Gigapath | GasHisDB | 0.995 (0.994-0.997) | 0.996 (0.994-0.997) | 1.000 (1.000-1.000) |
| GPFM | GasHisDB | 0.997 (0.996-0.998) | 0.997 (0.996-0.998) | 1.000 (1.000-1.000) |
| ResNet50 | Center-3-GC* | 0.551 (0.539-0.563) | 0.525 (0.506-0.547) | 0.721 (0.696-0.742) |
| Phikon | Center-3-GC* | 0.751 (0.730-0.769) | 0.729 (0.711-0.746) | 0.812 (0.789-0.834) |
| Ctranspath | Center-3-GC* | 0.708 (0.689-0.727) | 0.716 (0.696-0.736) | 0.777 (0.753-0.799) |
| UNI | Center-3-GC* | 0.819 (0.800-0.835) | 0.841 (0.823-0.857) | 0.796 (0.772-0.819) |
| CONCH | Center-3-GC* | 0.623 (0.605-0.641) | 0.634 (0.614-0.654) | 0.763 (0.741-0.785) |
| PLIP | Center-3-GC* | 0.555 (0.534-0.576) | 0.505 (0.486-0.525) | 0.598 (0.572-0.622) |
| CHIEF | Center-3-GC* | 0.681 (0.660-0.701) | 0.692 (0.673-0.713) | 0.746 (0.723-0.768) |
| Prov-Gigapath | Center-3-GC* | 0.819 (0.800-0.836) | 0.841 (0.825-0.859) | 0.794 (0.772-0.820) |
| GPFM | Center-3-GC* | 0.852 (0.837-0.870) | 0.886 (0.871-0.900) | 0.828 (0.804-0.851) |
@@
| ResNet50 | 28.17%(26.63%-29.70%) | 86.52%(85.43%-87.61%) | 57.32% (56.41%-58.28%) |
| Phikon | 30.78%(29.28%-32.29%) | 87.20%(86.13%-88.27%) | 58.97% (58.10%-59.93%) |
| Ctranspath | 31.11%(29.58%-32.65%) | 87.51%(86.44%-88.58%) | 59.35% (58.42%-60.28%) |
| UNI | 33.85%(32.28%-35.42%) | 88.69%(87.64%-89.74%) | 61.28% (60.39%-62.23%) |
| CONCH | 37.08%(35.40%-38.77%) | 88.51%(87.49%-89.53%) | 62.84% (61.84%-63.81%) |
| PLIP | 30.83%(29.29%-32.37%) | 88.02%(86.94%-89.09%) | 59.42% (58.48%-60.35%) |
| CHIEF | 32.11% (30.49%-33.60%) | 88.36% (87.28%-89.46%) | 60.23% (59.26%-61.18%) |
| Prov-Gigapath | 33.46% (31.80%-35.04%) | 88.35% (87.26%-89.40%) | 60.91% (59.88%-61.90%) |
| GPFM | 34.26%(32.67%-35.84%) | 88.41%(87.32%-89.49%) | 61.39% (60.39%-62.30%) |
| BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE-L | CE ACC | |
| ResNet50 | 0.386 | 0.324 | 0.301 | 0.157 | 0.230 | 0.456 | 0.482 |
| Phikon | 0.359 | 0.323 | 0.322 | 0.189 | 0.227 | 0.450 | 0.536 |
| Ctranspath | 0.386 | 0.333 | 0.316 | 0.162 | 0.238 | 0.459 | 0.462 |
| UNI | 0.381 | 0.322 | 0.315 | 0.202 | 0.231 | 0.458 | 0.482 |
| CONCH | 0.386 | 0.332 | 0.314 | 0.177 | 0.234 | 0.456 | 0.487 |
| PLIP | 0.388 | 0.317 | 0.288 | 0.148 | 0.225 | 0.457 | 0.474 |
| CHIEF | 0.400 | 0.350 | 0.335 | 0.206 | 0.245 | 0.474 | 0.497 |
| Prov-Gigapath | 0.381 | 0.322 | 0.303 | 0.179 | 0.234 | 0.470 | 0.526 |
| GPFM | 0.395 | 0.345 | 0.326 | 0.214 | 0.240 | 0.470 | 0.503 |
@@
| BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE-L | |
| Breast | ||||||
| ResNet50 | 0.228±0.007 | 0.079±0.004 | 0.034±0.003 | 0.016±0.002 | 0.081±0.002 | 0.157±0.004 |
| Phikon | 0.416±0.006 | 0.312±0.005 | 0.251±0.004 | 0.208±0.004 | 0.194±0.003 | 0.364±0.005 |
| Ctranspath | 0.254±0.007 | 0.131±0.004 | 0.078±0.002 | 0.047±0.002 | 0.097±0.004 | 0.192±0.004 |
| UNI | 0.361±0.006 | 0.255±0.005 | 0.198±0.004 | 0.161±0.003 | 0.162±0.004 | 0.306±0.005 |
| CONCH | 0.265±0.006 | 0.163±0.005 | 0.113±0.004 | 0.084±0.003 | 0.117±0.003 | 0.226±0.005 |
| PLIP | 0.269±0.005 | 0.148±0.003 | 0.093±0.003 | 0.061±0.002 | 0.106±0.003 | 0.201±0.002 |
| CHIEF | 0.272±0.010 | 0.151±0.006 | 0.093±0.004 | 0.060±0.003 | 0.117±0.004 | 0.209±0.005 |
| Prov-Gigapath | 0.334±0.007 | 0.230±0.004 | 0.175±0.005 | 0.139±0.003 | 0.148±0.004 | 0.278±0.005 |
| GPFM | 0.390±0.007 | 0.289±0.004 | 0.231±0.005 | 0.192±0.004 | 0.182±0.003 | 0.346±0.006 |
| Lung | ||||||
| ResNet50 | 0.224±0.008 | 0.085±0.005 | 0.035±0.003 | 0.016±0.002 | 0.078±0.002 | 0.159±0.004 |
| Phikon | 0.405±0.009 | 0.284±0.006 | 0.211±0.005 | 0.162±0.004 | 0.179±0.004 | 0.346±0.007 |
| Ctranspath | 0.150±0.011 | 0.066±0.006 | 0.032±0.004 | 0.015±0.002 | 0.058±0.004 | 0.131±0.006 |
| UNI | 0.329±0.009 | 0.220±0.006 | 0.158±0.004 | 0.119±0.003 | 0.140±0.004 | 0.279±0.006 |
| CONCH | 0.229±0.008 | 0.134±0.005 | 0.088±0.004 | 0.061±0.003 | 0.091±0.003 | 0.188±0.005 |
| PLIP | 0.198±0.007 | 0.079±0.005 | 0.035±0.004 | 0.014±0.003 | 0.072±0.003 | 0.139±0.004 |
| CHIEF | 0.209±0.012 | 0.098±0.007 | 0.044±0.004 | 0.019±0.003 | 0.079±0.004 | 0.167±0.007 |
| Prov-Gigapath | 0.308±0.009 | 0.199±0.006 | 0.140±0.004 | 0.102±0.003 | 0.132±0.004 | 0.260±0.006 |
| GPFM | 0.349±0.008 | 0.235±0.006 | 0.173±0.005 | 0.132±0.004 | 0.152±0.004 | 0.300±0.007 |
| Kidney | ||||||
| ResNet50 | 0.426±0.006 | 0.281±0.004 | 0.202±0.003 | 0.153±0.003 | 0.187±0.002 | 0.320±0.004 |
| Phikon | 0.500±0.007 | 0.375±0.006 | 0.300±0.005 | 0.247±0.005 | 0.225±0.004 | 0.406±0.005 |
| Ctranspath | 0.415±0.011 | 0.269±0.007 | 0.193±0.005 | 0.147±0.004 | 0.184±0.005 | 0.318±0.007 |
| UNI | 0.450±0.006 | 0.333±0.005 | 0.267±0.005 | 0.222±0.004 | 0.201±0.004 | 0.364±0.005 |
| CONCH | 0.420±0.006 | 0.280±0.005 | 0.203±0.004 | 0.156±0.004 | 0.185±0.003 | 0.318±0.004 |
| PLIP | 0.400±0.006 | 0.259±0.004 | 0.185±0.003 | 0.141±0.002 | 0.171±0.002 | 0.303±0.003 |
| CHIEF | 0.384±0.004 | 0.233±0.005 | 0.153±0.004 | 0.106±0.003 | 0.153±0.004 | 0.280±0.005 |
| Prov-Gigapath | 0.416±0.006 | 0.292±0.005 | 0.224±0.005 | 0.179±0.004 | 0.184±0.004 | 0.329±0.005 |
| GPFM | 0.504±0.008 | 0.381±0.006 | 0.307±0.005 | 0.255±0.004 | 0.226±0.004 | 0.407±0.005 |
| BLEU-1 | BLEU-2 | BLEU-3 | BLEU-4 | METEOR | ROUGE-L | |
| ResNet50 | 0.596±0.019 | 0.496±0.021 | 0.424±0.022 | 0.369±0.023 | 0.301±0.011 | 0.564±0.021 |
| Phikon | 0.655±0.025 | 0.577±0.027 | 0.522±0.029 | 0.481±0.030 | 0.347±0.015 | 0.623±0.026 |
| Ctranspath | 0.643±0.020 | 0.556±0.022 | 0.495±0.023 | 0.450±0.025 | 0.334±0.012 | 0.598±0.022 |
| UNI | 0.609±0.023 | 0.533±0.025 | 0.482±0.027 | 0.444±0.028 | 0.327±0.014 | 0.596±0.024 |
| CONCH | 0.641±0.019 | 0.555±0.023 | 0.495±0.025 | 0.450±0.026 | 0.337±0.013 | 0.599±0.022 |
| PLIP | 0.650±0.021 | 0.560±0.024 | 0.498±0.026 | 0.451±0.027 | 0.338±0.013 | 0.599±0.023 |
| CHIEF | 0.594±0.021 | 0.494±0.023 | 0.425±0.025 | 0.372±0.026 | 0.298±0.013 | 0.561±0.025 |
| Prov-Gigapath | 0.637±0.022 | 0.555±0.025 | 0.497±0.026 | 0.454±0.027 | 0.338±0.013 | 0.601±0.023 |
| GPFM | 0.651±0.021 | 0.569±0.023 | 0.512±0.025 | 0.470±0.026 | 0.343±0.013 | 0.606±0.025 |
| Score: 0 | Score: 0.3 | Score: 0.7 | Score: 1 | Avg. | |
| Breast | |||||
| ResNet50 | 184 | 4 | 0 | 0 | 0.01 |
| Phikon | 18 | 136 | 34 | 0 | 0.34 |
| Ctranspath | 80 | 88 | 20 | 0 | 0.21 |
| UNI | 9 | 134 | 45 | 0 | 0.38 |
| CONCH | 36 | 125 | 27 | 0 | 0.30 |
| PLIP | 118 | 69 | 1 | 0 | 0.11 |
| CHIEF | 25 | 139 | 24 | 0 | 0.31 |
| Prov-Gigapath | 24 | 127 | 37 | 0 | 0.34 |
| GPFM | 15 | 126 | 47 | 0 | 0.38 |
| Lung | |||||
| ResNet50 | 153 | 3 | 1 | 0 | 0.01 |
| Phikon | 7 | 109 | 41 | 0 | 0.39 |
| Ctranspath | 19 | 87 | 51 | 0 | 0.39 |
| UNI | 13 | 109 | 35 | 0 | 0.36 |
| CONCH | 11 | 120 | 26 | 0 | 0.35 |
| PLIP | 53 | 104 | 0 | 0 | 0.20 |
| CHIEF | 15 | 113 | 29 | 0 | 0.35 |
| Prov-Gigapath | 13 | 111 | 33 | 0 | 0.36 |
| GPFM | 6 | 100 | 51 | 0 | 0.42 |
| Kidney | |||||
| ResNet50 | 175 | 0 | 0 | 0 | 0.00 |
| Phikon | 5 | 86 | 84 | 0 | 0.48 |
| Ctranspath | 32 | 76 | 67 | 0 | 0.39 |
| UNI | 2 | 91 | 82 | 0 | 0.48 |
| CONCH | 11 | 118 | 46 | 0 | 0.39 |
| PLIP | 172 | 3 | 0 | 0 | 0.01 |
| CHIEF | 9 | 100 | 66 | 0 | 0.44 |
| Prov-Gigapath | 5 | 86 | 82 | 2 | 0.49 |
| GPFM | 2 | 83 | 90 | 0 | 0.50 |
@@
| CHIEF [64] | Public+Private | 60K | 15M | SwinTrans. | 28M | MoCoV3+CLIP |
| Prov-Gigapath [56] | Private | 171K | 1.3B | Vit-g | 1.1B | DINOv2+MAE |
| GPFM (our) | 33 Public datasets | 72K | 190M | ViT-L | 307M | UDK |